# Scalable Bayesian Learning with posteriors

### v1

@@ 4 Stochastic gradient descent, SGMCMC and deep ensembles
This section introduces a tempered version of SGMCMC, as implemented in posteriors, and demonstrates how this provides a unified transition between stochastic gradient descent (SGD) and SGMCMC methods. We then show how this points to a small modification of deep ensembles [43] to ensure convergence to the posterior. Comprehensive details on SGMCMC are found in Appendix A.3.
@@ 4.1 A unified framework for SGD and SGMCMC
Consider targeting a tempered target distribution $\pi_{\operatorname{\mathcal{T}}}(z)\propto\pi(z)^{\frac{1}{N\operatorname{\mathcal{T}}}}$ for temperature $\operatorname{\mathcal{T}}\in[0,\infty)$, which can be considered as a synthetic hyperparameter controlling the variance of $\pi_{\operatorname{\mathcal{T}}}(z)$ [77, 18]. Here $z=(\theta,\omega)$ where $\omega$ are any auxiliary parameters such as momenta [12] or thermostat [18]. The distribution $\pi(z)$ is thus an extended target with $p(\theta\mid y_{1:N})$ marginal in $\theta$.
An SDE is invariant for $\pi_{\operatorname{\mathcal{T}}}(z)$ if and only if it has the following form [48]
@@
where the second line comes from rescaling $\operatorname{\mathrm{D}}(z),\operatorname{\mathrm{Q}}(z)\to\operatorname{\mathcal{T}}\operatorname{\mathrm{D}}(z),\operatorname{\mathcal{T}}\operatorname{\mathrm{Q}}(z)$  [^1] Equivalently we can rescale $dt\to\operatorname{\mathcal{T}}dt$ as Brownian motion $dw$ scales with $\sqrt{t}$, see (1.27) in [63].
@@
Intuitively, the low-temperature limit $\operatorname{\mathcal{T}}\to 0$ converts the target distribution $\pi(z)$ into one with all probability mass concentrated on (local) maxima and transparently the SDE (3) into an ODE
@@
In Appendix C, we give specific examples demonstrating the tempered equivalence between vanilla SGD and stochastic gradient Langevin dynamics (SGLD) [76] as well as SGD with momenta and stochastic gradient Hamiltonian Monte Carlo (SGHMC) [12].
@@ 4.2 Parallel SGMCMC (as Bayesian deep ensembles)
Deep ensembles [43] represent a simple to implement and parallelizable route to uncertainty quantification via multiple runs of gradient descent with different random seeds (for initialization and batch shuffling).
Thus, the arguments above point to a formulation of Bayesian deep ensembles by converting the gradient descent SDE (4) typically used for deep ensemble training into a Bayesian version that is stationary for the posterior distribution by setting $\operatorname{\mathcal{T}}=N^{-1}$, i.e. adding back the noise term in (3). Assuming $\operatorname{\mathrm{D}}$ and $\operatorname{\mathrm{Q}}$ are chosen to be independent of $z$ this becomes
@@
The core difference between SGMCMC and the Bayesian deep ensemble or parallel SGMCMC implementation above is that in SGMCMC, samples are collected from a single trajectory, whereas in parallel SGMCMC, multiple trajectories are simulated with only the last sample being retained (although multiple samples per trajectory can also be saved [53]). Parallel SGMCMC provides an asymptotically unbiased approximation to $\pi(z)$ and therefore $p(\theta~{}\mid~{}y_{1:N})$ as long as the criteria discussed in A.3 are met for a single chain with the main one being the decaying learning rate which we are more accepting of in the parallel regime since the exploration requirements are divided across trajectories rather relying on a single trajectory. Parallel MCMC is also a bright avenue of research for full-batch inference [53, 14], and we expect these techniques to become even more prevalent with further advances in parallel hardware.
We note that in the large data regime, the additional noise term in (5) versus (4) becomes very small, and so existing deep ensembles [43] can already be seen as approximating the posterior in some sense (with a small bias).
Figure 3 visualizes the behaviour of the techniques discussed in this section for a two-dimensional multi-modal double well posterior. We see clearly that although a deep ensemble finds all modes, it offers no within-mode uncertainty. In contrast, serial SGMCMC explores well within modes but struggles to transition between modes. Parallel SGMCMC combines the benefits of both methods to provide a more representative posterior approximation. In more realistic high dimensional, large $N$ settings runtime and memory requirements become a key consideration. Serial SGMCMC requires less total computation since it runs only a single trajectory. However, the trajectory must be significantly longer and cannot be parallelized; therefore, serial SGMCMC is perhaps not as well suited to modern computational architectures as the deep ensemble or parallel SGMCMC approaches.
@@ 5 Experiments
We now present some experiments with posteriors that demonstrate the three key benefits of the Bayesian paradigm for machine learning as discussed in Section 1. Reproducible code for all experiments can be found in the examples folder of the posteriors repository.
@@ 5.1 Cold Posterior Effect
The purpose of this experiment is to investigate the generalization capabilities of the Bayesian methods discussed in 2. We use posteriors’ unified framework, Figure 3, to easily swap between Bayesian algorithms. Of particular note, we observe that the cold posterior effect is very prominent in Gaussian approximations whilst being mild for SGMCMC which can significantly outperform gradient descent approaches. Extensive experimental details can be found in Appendix D.
The cold posterior effect [78, 77, 2, 35] is a phenomenon that has been observed for approximate Bayesian machine learning models where increased predictive performance has been observed when targeting $p_{T}(\theta\mid y_{1:N})\propto p(\theta\mid y_{1:N})^{\frac{1}{T}}$ (as discussed in 4 with $\operatorname{\mathcal{T}}=TN^{-1}$) for $T<1$, which appears at odds with the Bayesian paradigm [2].
Figure 4: Investigation into the cold-posterior effect for a range of posteriors algorithms for a CNN-LSTM model [77] on the IMDB dataset [49]. Non-linearized Laplace EF and Laplace GGN are indistinguishable. All Gaussian approximations are diagonal. All approaches display error bars with one standard deviation over 5 random seeds.
Following [77, 35], we train a CNN-LSTM model with 2.7 million parameters on the IMDB [49] for binary classification of positive or negative reviews. As in [35], we use a diagonal Gaussian prior with variance $1/40$ for all methods.
@@
The Bayesian methods, particularly VI and SGHMC can significantly outperform gradient descent for the MAP when evaluating on the test data. A simple Laplace approximation performs badly, however, improved performance is gained via the linearzation discussed in 2. Laplace GGN, which is equivalent to $F_{C}$ in this case (and in most machine learning models, see Appendix A.1), improves over $F_{E}$, which is in agreement with [42]. We observe that the reverse is true for VI and that the linearization is actually detrimental. We posit that this is due to the VI optimization acting in parameter space without knowledge of forward linearization. We observe a strong cold posterior effect for the Gaussian approaches with significantly superior performance achieved at $T<1$. However, for the (non-Gaussian) SGMCMC approaches, only a very mild cold posterior approach is observed. This is evidence in agreement with [35] that the cold posterior effect is more a result of the crudeness of posterior approximations than the of Bayesian learning itself (1). Finally, we observe that the parallel SGHMC approach (Bayesian deep ensemble) slightly outperforms the single trajectory serial approach, relatably Figure 3. And further, reducing the temperature deteriorates performance (observing from Section 4 and Appendix C that traditional deep ensemble is recovered as parallel SGHMC with $T=0$).
@@ 5.2 Continual learning with LoRA
The purpose of this experiment is to apply a simple Bayesian technique, in particular, an empirical Fisher Laplace approximation to mitigate catastrophic forgetting in an online learning problem. It also demonstrates the composability of posteriors with parameter-efficient fine-tuning (PEFT) [52] and a large language model in Llama 2 [74] from the PyTorch [62] ecosystem. Extensive experimental details can be found in Appendix E.
We consider the problem of episodal continual learning [40, 59] where each episode’s data, $y_{n}$, is one of 20 books from the PG19 [65] dataset with 15% of each book held out for testing. We compare simple online gradient descent with an approximation to the Bayesian online update (1) $p(\theta\mid y_{1:n})\propto p(\theta\mid y_{1:n-1})p(y_{n}\mid\theta)$ in Figure 6 and 6. Our baseline uses AdamW [47] with a constant learning rate, and we reset all hyperparameters (including momentum) between each episode. The Laplace posterior from the previous episode is then used as the prior for the next episode following the suggested amendment to the EWC method [40] recommended in [32]. We also train an offline version of the same model with access to all books to indicate the model’s total capacity (although unrealistic of an online pipeline).
@@
We use LoRA [31] to fine-tune the query, key and output weight matrices in the last attention layer of the 7B parameter model Llama 2 model [74]. LoRA computes a low-rank approximation to the weight updates and is implemented in the PEFT library [52], which integrates easily into posteriors code. We use standard choices for the rank ($r=8$) and scaling ($\alpha=32$) parameters.
Figure 6 shows test loss for each episode, over all four episodes. The Laplace method maintains low loss in early tasks, while SGD forgets. For example, in the top row, the Laplace approximation stays low throughout training demonstrating that it continues to perform well on task 0 even though it is now being trained on data from tasks 1-3. In contrast, continually applying gradient descent quickly decreases the model’s performance on task 0. Figure 6 shows the difference of test loss from baseline, averaged over all episodes seen thus far. We see that the Laplace approximation approach maintains low loss averaged over all tasks.
@@ 5.3 Bayesian Llama 3
The purpose of this experiment is to show how we can use posteriors to fine-tune a very recent large language model to provide the capability to decompose uncertainty (2) and provide improved out-of-distribution detection. Extensive experimental details can be found in Appendix F.
We fine-tune the last attention layer of the very recent 8B Llama 3 model [1] (resulting in 218 million trainable parameters) on the TQA [37] dataset, which consists of scientific textbooks. We fine-tune a baseline with SGD and a size 10 ensemble trained in parallel with posteriors’ SGHMC. We also compare against the base Llama 3 model without fine-tuning. We then evaluate the models by asking them to complete 100 scientific statements (in-distribution) with the last word removed. We then repeat the experiment with the statements translated into Samoan (an out-of-distribution language for TQA [37]).
Figure 7: Distributions of predictive uncertainties with Llama 3.
We can see from Figure 7 that although total uncertainty provides some indication for classifying the Samoan text as out-of-distribution, the epistemic uncertainty of parallel SGMHC offers a clearer split. Table 1 validates this with the AUROC metric [41], whilst also displaying the categorical loss values for the next token of the English statements to verify a successful fine-tune.
Indeed, these are important results for mission-critical tasks. Expanding upon these techniques can result in higher reliability of language models as they can inform the user when the model feels high uncertain about its provided answers.
@@ 6 Related work
Section 2 (and more thoroughly Appendix A) summarizes much of the related work regarding Bayesian methodology. Additional comprehensive reviews can be found in [75] with [24] focussing on prior specifications (which we do not go into detail here but is handled with broad flexibility in posteriors) and [35] providing extensive experiments with full-batch MCMC to give a closer indication of the behaviour of the true Bayesian posterior.
There is a broad selection of software for uncertainty quantification; however, to our knowledge and experience, none satisfy all five points in 3. Stan [11] is widely used within the field of Bayesian statistics however does not support minibatching and is not easily extensible. Pyro [6] is a broad framework that supports a range of models and inference techniques; however, it is more heavyweight than posteriors and harder to compose with existing PyTorch code. BlackJAX [10] (and optax [16]) heavily inspired the posteriors functional approach. However, BlackJAX also has a primary focus on full-batch inference whilst posteriors is minibatch-first and, like other JAX [64, 17, 19] packages, does not compose with the PyTorch ecosystem.

### v2

@@ 4 Stochastic gradient descent, SGMCMC and deep ensembles
This section describes a tempered version of SGMCMC that introduces an additional temperature parameter that allows us to view SGMCMC as a generalization of stochastic gradient descent (SGD) and variants. This viewpoint has guided the implementation of posteriors allowing users access to SGMCMC within a familiar functional framework of popular optimization libraries such as torchopt (Ren et al., 2023) and optax (DeepMind et al., 2020), thus making Bayesian learning scalable and maximally accessible to the broader machine learning community. In the second half of this section, we show how this points to a small modification of deep ensembles (Lakshminarayanan et al., 2017) to ensure convergence to the posterior, and therefore how posteriors can be used to easily convert a deep ensemble pipeline into one of (parallel) Bayesian learning. Comprehensive details on SGMCMC are found in Appendix A.3.
The tempered connection between SGMCMC and SGD has been discussed in previous works for specific optimizers/samplers e.g. Wenzel et al. (2020), however here we give a complete characterization of the connection based on the framework in Ma et al. (2015) which dictates the implementations in posteriors. Additionally, to our knowledge, the corollary unifying deep ensembles and parallel SGMCMC is novel.
The minibatch-first nature of posteriors alongside the reframing of stochastic gradient optimization as SGMCMC results in an implementation with access to both approaches within a unified syntax as well as enabling the codesign of new algorithms, all within a familiar optimization-style framework.
@@ 4.1 A unified framework for SGD and SGMCMC
Consider targeting a tempered target distribution $\pi_{\operatorname{\mathcal{T}}}(z)\propto\pi(z)^{\frac{1}{N\operatorname{\mathcal{T}}}}$ for temperature $\operatorname{\mathcal{T}}\in[0,\infty)$, which can be considered as a synthetic hyperparameter controlling the variance of $\pi_{\operatorname{\mathcal{T}}}(z)$ (Wenzel et al., 2020; Ding et al., 2014). Here $z=(\theta,\omega)$ where $\omega$ are any auxiliary parameters such as momenta (Chen et al., 2014) or thermostat (Ding et al., 2014). The distribution $\pi(z)$ is thus an extended target with $p(\theta\mid y_{1:N})$ marginal in $\theta$.
An SDE is invariant for $\pi_{\operatorname{\mathcal{T}}}(z)$ if and only if it has the following form (Ma et al., 2015)
@@
where the second line comes from rescaling $\operatorname{\mathrm{D}}(z),\operatorname{\mathrm{Q}}(z)\to\operatorname{\mathcal{T}}\operatorname{\mathrm{D}}(z),\operatorname{\mathcal{T}}\operatorname{\mathrm{Q}}(z)$  [^1] Equivalently we can rescale $dt\to\operatorname{\mathcal{T}}dt$ as Brownian motion ($dw$) scales with $\sqrt{t}$, see (1.27) in Pavliotis (2014).
@@
Intuitively, the low-temperature limit $\operatorname{\mathcal{T}}\to 0$ converts the target distribution $\pi(z)$ into one with all probability mass concentrated on (local) maxima and transparently the SDE equation 3 into an ODE
@@
All SGMCMC algorithms in posteriors feature the temperature $\mathcal{T}$ as a parameter with $\mathcal{T}=0$ resulting in a valid SGD variant (specific examples can be found in Appendix C). The extensibility and strict support for minibatches allow a unified posteriors’ implementation and new procedures immediately apply to both SGMCMC and SGD (e.g. step-size adaptation, which has in classical Bayesian statistics utilized Metropolis acceptance rates which do not apply to SGMCMC).
@@ 4.2 Parallel SGMCMC (as Bayesian deep ensembles)
Deep ensembles (Lakshminarayanan et al., 2017) represent a simple to implement and parallelizable route to uncertainty quantification via multiple runs of gradient descent with different random seeds (for initialization and batch shuffling). Thus, the arguments above point to the formulation of Bayesian deep ensembles, enabled by posteriors, that converts the gradient descent ODE equation 4 typically used for deep ensemble training into a Bayesian version that is stationary for the posterior distribution by setting $\operatorname{\mathcal{T}}=N^{-1}$, i.e. adding back the noise term in equation 3. Assuming $\operatorname{\mathrm{D}}$ and $\operatorname{\mathrm{Q}}$ are chosen to be independent of $z$ this becomes
@@
The core difference between SGMCMC and the Bayesian deep ensemble or parallel SGMCMC implementation above is that in SGMCMC, samples are collected from a single trajectory, whereas in parallel SGMCMC, multiple trajectories are simulated with only the last sample being retained (although multiple samples per trajectory can also be saved (Margossian et al., 2021)). Parallel SGMCMC provides an asymptotically unbiased approximation to $\pi(z)$ and therefore $p(\theta~{}\mid~{}y_{1:N})$ as long as the criteria discussed in Appendix A.3 are met for a single chain.
We note that in the large data regime, the additional noise term in equation 5 versus equation 4 becomes very small, and so existing deep ensembles (Lakshminarayanan et al., 2017) can already be seen as approximating the posterior in some sense (with a small bias).
The functional framework of posteriors (in combination with torch.func) allows easy parallelization of the above unified SGMCMC and SGD algorithms into a parallel SGMCMC implementation that generalizes deep ensembles with minimal code change.
Figure 3 visualizes the behaviour of the techniques discussed in this section for a two-dimensional multi-modal double well posterior. We see clearly that although a deep ensemble finds all modes, it offers no within-mode uncertainty. In contrast, serial SGMCMC explores well within modes but struggles to transition between modes. Parallel SGMCMC combines the benefits of both methods to provide a more representative posterior approximation. In more realistic high dimensional, large $N$ settings runtime and memory requirements become a key consideration. Serial SGMCMC requires less total computation since it runs only a single trajectory. However, the trajectory must be significantly longer and cannot be parallelized; therefore, serial SGMCMC is perhaps not as well suited to modern computational architectures as the deep ensemble or parallel SGMCMC approaches. However, parallel SGMCMC naturally shares many of the caveats of deep ensembles. Notably, the resource requirements are significant to parallelize trajectories, although we note the opportunity to collect multiple samples per trajectory (Margossian et al., 2021). Another point of note is that communication costs between cores are expensive and hyperparameter tuning may require alternative approaches to the serial setting and typically need more care (than optimization) to preserve the correct target distribution across trajectories.
@@ 5 Experiments
We now present experiments investigating the Bayesian learning benefits motivated in Section 1. The functional and swappable nature of posteriors makes these experiments particularly accessible across a range of Bayesian learning algorithms. Notably, its scalability and composability with the PyTorch ecosystem allow us to apply Bayesian learning to a host of pre-trained HuggingFace models (Wolf et al., 2020). Reproducible code for all experiments (and more) can be found in the examples folder of the posteriors repository.
@@ 5.1 Cold Posterior Effect
This experiment fully utilizes posteriors swappability between algorithms to investigate the generalization capabilities of Bayesian learning. As seen in Figure 3, within a posteriors implementation the majority of the code is algorithm agnostic and we can easily swap approaches by only changing the build of the transform object and its configuration parameters. Extensive experimental details can be found in Appendix D.
The cold posterior effect (Wilson and Izmailov, 2020; Wenzel et al., 2020; Aitchison, 2020; Izmailov et al., 2021) is a phenomenon that has been observed for approximate Bayesian machine learning models where increased predictive performance has been observed when targeting $p_{T}(\theta\mid y_{1:N})\propto p(\theta\mid y_{1:N})^{\frac{1}{T}}$ (as discussed in Section 4 with $\operatorname{\mathcal{T}}=TN^{-1}$) for $T<1$, which appears at odds with the Bayesian paradigm (Aitchison, 2020).
Figure 4: Investigation into the cold-posterior effect for a range of posteriors algorithms for a CNN-LSTM model (Wenzel et al., 2020) on the IMDB dataset (Maas et al., 2011). Non-linearized Laplace EF and Laplace GGN are indistinguishable (left panel). All Gaussian approximations are diagonal. All approaches display error bars with one standard deviation over 5 random seeds.
Following Wenzel et al. (2020); Izmailov et al. (2021), we train a CNN-LSTM model with 2.7 million parameters on the IMDB (Maas et al., 2011) for binary classification of positive or negative reviews. As in Izmailov et al. (2021), we use a diagonal Gaussian prior with variance $1/40$ for all methods.
@@
The Bayesian methods, particularly VI and SGHMC can significantly outperform gradient descent for the MAP when evaluating on the test data. A simple Laplace approximation performs badly, however, improved performance is gained via the linearization discussed in Section 2. Laplace GGN, which is equivalent to $F_{C}$ in this case (and in most machine learning models, see Appendix A.1), improves over $F_{E}$, which is in agreement with Kunstner et al. (2019). We observe that the reverse is true for VI and that the linearization is actually detrimental. We posit that this is due to the VI optimization acting in parameter space without knowledge of forward linearization. We observe a strong cold posterior effect for the Gaussian approaches with significantly superior performance achieved at $T<1$. However, for the (non-Gaussian) SGMCMC approaches, only a very mild cold posterior approach is observed. This is evidence in agreement with Izmailov et al. (2021) that the cold posterior effect is more a result of the crudeness of posterior approximations than the of Bayesian learning itself equation 1. Finally, we observe that the parallel SGHMC approach (Bayesian deep ensemble) slightly outperforms the single trajectory serial approach, relatably Figure 3. And further, reducing the temperature deteriorates performance (observing from Section 4 and Appendix C that traditional deep ensemble is recovered as parallel SGHMC with $T=0$).
@@ 5.2 Continual learning with LoRA
The purpose of this experiment is to apply a simple Bayesian technique made accessible by posteriors, namely, an empirical Fisher Laplace approximation, to mitigate catastrophic forgetting in an online setting with a large-scale model. Here, the composability of posteriors, as well as the flexibility of the functional framework, is vital to make use of parameter-efficient fine-tuning (PEFT) (Mangrulkar et al., 2022) and the Llama 2 (Touvron et al., 2023) large language model from the PyTorch (Paszke et al., 2019) ecosystem. Extensive details can be found in Appendix E.
@@
We consider the problem of episodal continual learning (Kirkpatrick et al., 2017; Nguyen et al., 2017) where each episode’s data, $y_{n}$, is one of 20 books from the PG19 (Rae et al., 2019) dataset with 15% of each book held out for testing. We compare simple online gradient descent with an approximation to the Bayesian online update equation 1 $p(\theta\mid y_{1:n})\propto p(\theta\mid y_{1:n-1})p(y_{n}\mid\theta)$ in Figures 6 and 6. Our baseline implements AdamW via torchopt’s integration within the posteriors unified API. The Laplace posterior from the previous episode is then used as the prior for the next episode following the suggested amendment to the EWC method (Kirkpatrick et al., 2017) recommended in (Huszár, 2018). We also train an offline version of the same model with access to all books to indicate the model’s total capacity (although unrealistic of an online pipeline).
We use LoRA (Hu et al., 2021) to fine-tune the query, key and output weight matrices in the last attention layer of the 7B parameter model Llama 2 model (Touvron et al., 2023). LoRA computes a low-rank approximation to the weight updates and is implemented in the PEFT library Mangrulkar et al. (2022), which integrates easily into functional posteriors code. We use standard choices for the rank ($r=8$) and scaling ($\alpha=32$) parameters.
Figure 6 shows test loss for each episode, over all four episodes. The Laplace method maintains low loss in early tasks, while SGD forgets. For example, in the top row, the Laplace approximation stays low throughout training demonstrating that it continues to perform well on task 1 even though it is now being trained on data from tasks 2-4. In contrast, continually applying gradient descent quickly decreases the model’s performance on task 1. Figure 6 shows the difference of test loss from baseline, averaged over all episodes seen thus far. We see that the Laplace approximation approach maintains low loss averaged over all tasks.
@@ 5.3 Bayesian Llama 3
Figure 7: Distributions of predictive uncertainties with Llama 3.
We now show how posteriors can be used to fine-tune a very recent large language model to provide the capability to decompose uncertainty (equation 2) and provide improved out-of-distribution detection. The flexible and scalable design of posteriors allows us to easily use subspace methods for fine-tuning a large pre-trained model in parallel using the novel Bayesian generalization of deep ensembles introduced in Section 4.2. Additionally, posteriors’ functional approach allows us to easily and efficiently map over the ensemble in forward calls at inference time. Extensive experimental details can be found in Appendix F.
We fine-tune the last attention layer of the 8B Llama 3 model (AI@Meta, 2024) (resulting in 218 million trainable parameters) on the TQA (Kembhavi et al., 2017) dataset, which consists of scientific textbooks. We fine-tune a baseline with SGD and a size 10 ensemble trained in parallel with posteriors’ SGHMC. We also compare against the base Llama 3 model without fine-tuning. We evaluate the models by asking them to complete 100 scientific statements (in-distribution) with the last word removed. We then repeat the experiment with the statements translated into Samoan, an out-of-distribution language for TQA (Kembhavi et al., 2017).
We can see from Figure 7 that although total uncertainty provides some indication for classifying the Samoan text as out-of-distribution, the epistemic uncertainty of parallel SGMHC offers a clearer split. Table 1 validates this with the AUROC metric (Kuhn et al., 2023), whilst also displaying the categorical loss values for the next token of the English statements to verify a successful fine-tune. Overall, we can see a significant advantage in out-of-distribution detection via the Bayesian decomposition of uncertainty, whereas the point-estimate-based methods (MLE/MAP) cannot decompose uncertainty to differentiate epistemic uncertainty from aleatoric or semantic uncertainty which is key for open-ended natural language generation (Kuhn et al., 2023).
@@ 6 Related work
Section 2 (and more thoroughly Appendix A) summarizes much of the related work regarding Bayesian methodology. Additional comprehensive reviews can be found in Wang and Yeung (2020) with Fortuin (2022) focussing on prior specifications (which we do not go into detail here but is handled with broad flexibility in posteriors) and Izmailov et al. (2021) providing extensive experiments with full-batch MCMC to give a closer indication of the behaviour of the true Bayesian posterior.
There is a broad selection of software for uncertainty quantification; however, to our knowledge and experience, none satisfy all five points in Section 3. Stan (Carpenter et al., 2017) is widely used within the field of Bayesian statistics however does not support minibatching and is not easily extensible. Pyro (Bingham et al., 2019) is a broad framework that supports a range of models and inference techniques; however, it is more heavyweight than posteriors and harder to compose with existing PyTorch code. BlackJAX (Cabezas et al., 2024) (and optax (DeepMind et al., 2020)) heavily inspired the posteriors functional approach. However, BlackJAX also has a primary focus on full-batch inference whilst posteriors is minibatch-first and, like other JAX (Phan et al., 2019; Detommaso et al., 2023; Duffield, 2021) packages, does not compose with the PyTorch ecosystem.